OpenStudy (anonymous):

A nice series problem, it may be helpful. Calculate the sum of: S[(2n+1)/(n+2)!] n=1 to infinity (An advice: use the serie for number e)

OpenStudy (anonymous):

\[\sum_{1}^{\infty} (2n+1)/(n+2)!\]

OpenStudy (anonymous):

I'm not asking, it's just for someone who wants to try

OpenStudy (zarkon):

pretty trivial...\[\frac{7}{2}-e\] I would write \(2n+1\) as \(2n+4-3=2(n+2)-3\)

OpenStudy (anonymous):

True